# Differentially Private SGD Without Clipping Bias: An Error-Feedback Approach

### v1

@@ 3.2 Theoretical analysis
Assume the problem satisfies Assumptions 3.5, and 3.6, given constant $C$, by fixing the clipping thresholds $0<C_{1}\leq C_{2}\leq C/B$, independent of $G,\sigma$, and assume $\frac{B}{N}\leq\frac{1}{5}$. Choose DP noise standard deviation $\sigma_{1}$ as
@@
Note that although Assumptions 3.5, and 3.6 are used in the proof, the result does not rely on the specific values of the bounds, which can be arbitrarily large. Due to the accumulated influence of the update of $\mathbf{e}^{t}$, the DiceSGD requires larger DP-noise than the DPSGD algorithm (larger by a constant multiplicative factor). The detailed proof is given in Appendix A.2. By optimizing $T$ we have the following utility-privacy trade-off for DiceSGD.
@@
Under the same assumptions of Theorem 3.7, choose the stepsize $\eta=\mathcal{O}(T^{-(1-\nu)}),$ where $0<\nu\leq\frac{1}{2}$ is a small constant, and clipping thresholds $0<C_{1}\leq C_{2}\leq C/B$, and choose noise multiplier $\sigma_{1}^{2}$ as Theorem 3.8. By running DiceSGD for $T=\mathcal{O}\left(\frac{N^{2}\epsilon^{2}}{\tilde{G}\log(1/\delta)}\right)$ iterations, the algorithm guarantees $(\epsilon,\delta)$-DP, while converging to a solution where the loss function satisfies:
$\operatorname{\mathbb{E}}[\mathcal{L}^{T+1}]=\mathcal{O}\left(\frac{\tilde{G}\log(N\epsilon)\log(1/\delta)}{N^{2}\epsilon^{2}}\right).$
A similar result can also be derived by replacing Assumption 3.3 with Assumption 3.4. The corollary indicates that when $N\rightarrow\infty$, the expected loss converges with rate $\mathcal{O}(\frac{\log(N)}{N^{2}})$ with arbitrary clipping thresholds $C_{2}\geq C_{1}>0$ and eliminates the constant clipping bias in DPSGD-GC.
@@ 4 Numerical experiments
We perform extensive evaluations of DiceSGD on image classification, and natural language processing (NLP) tasks to demonstrate its advantage over DPSGD-GC. The experiments were run on an Intel Xeon W-2102 CPU with an NVIDIA TITAN X GPU for image classification, and on an NVIDIA A100 GPU for NLP tasks. We conduct extra ablation studies on the choice of the clipping threshold $C_{1},C_{2}$ and learning rate $\eta$ on Cifar-10 and Cifar-100 datasets, which show that DiceSGD benefits from using a smaller clipping threshold and choosing $C_{2}=C_{1}$ gives the best result in most cases. More results and discussions are given in Appendix 4.1 due to the space limitation.
4.1 Image classification.
We use both Cifar-10 and Cifar-100 datasets for experiments and use ViT-small (Dosovitskiy et al., 2020) as the training model, which is pre-trained on Imagenet. In the experiment, we use the similar Adam variant of DPSGD-GC developed following Bu et al. (2021) to implement both DPSGD-GC and DiceSGD (See Appendix C.1 for details); we fine-tune the model for $3$ epochs with batch size $B=1000$ and ensure $(2,10^{-5})$-DP for both algorithms. The stepsize for DPSGD-GC and DiceSGD are selected through grid search from $\eta\in\{10^{-2},10^{-3},10^{-4}\}.$ The experiment results are shown in Table 3.
Table 3: Test accuracy of DPSGD-GC and DiceSGD on Cifar-10 and Cifar-100 datasets with different clipping thresholds.
@@
Ablation study
(a) Cifar-10
(b) Cifar-100
(a) Cifar-10
(b) Cifar-100
4.2 Natural language processing.
To validate the ability of DiceSGD for training larger models on different tasks, we further conduct experiments on the NLP task. Specifically, we fine-tune the GPT-2 model (Radford et al., 2018) on the E2E NLG Challenge for $10$ epochs with batch size $B=1000$, and report the standard metrics such as BLUE, ROUGE-L, etc., used in Hu et al. (2021) for evaluation. The results in Table 4 show that DiceSGD has better performance than DPSGD-GC.
@@ Appendix A Proof of results in Section 3
In this section, we provide detailed proofs of the theorems in Section 3. We find the following equations useful:
@@
Lemma A.2 (Theorem 2 Karimi et al. (2016a)).
Assume a function $f$ satisfies Assumptions 3.2 and 3.3 with constants $L$ and $\mu$, respectively, then we have $f$ satisfies Assumption 3.4 with parameter $\mu$.
We provide the statements of Theorem 3.7 under the PL condition assumption 3.4 and Theorem 3.8 below for convenience.
@@
Assume the problem satisfies Assumption 3.2, 3.4, 3.5, and 3.6. Given any constant DP noise multiplier $\sigma_{1}$, by running DiceSGD (Algorithm 2) for $T$ iterations, choosing stepsize $\eta=\min\{\frac{1}{2\mu},\frac{3}{32L},\frac{\sqrt{C_{2}}}{5L\sqrt{C_{2}+T\max\{0,G^{\prime}-C_{2}\}}}\}$, $G^{\prime}=\max\{0,G+\sigma-C_{1}\}$, clipping thresholds $C_{2}\geq C_{1}>0$. It satisfies
$\operatorname{\mathbb{E}}[\mathcal{L}^{T+1}]\leq(1-\mu\eta)^{T}\mathcal{L}^{0}+\frac{\eta L}{2\mu}\left(d\sigma_{1}^{2}+\frac{5\sigma^{2}}{4B}+\frac{C_{1}^{2}}{2}\right)$ (7)
where $\mathcal{L}^{t}=f(\tilde{\mathbf{x}}^{t})-f^{\star}+\frac{\eta^{3}L^{2}}{6}\left\lVert\mathbf{e}^{t}\right\rVert^{2}+\frac{\eta(1-L\eta)}{2}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}.$
We see that Theorem 3.7 can be derived from Theorem A.3 and Lemma A.2, because whenever we use Assumption 3.4 for proving Theorem A.3, we can apply Lemma A.2 and use Assumption 3.3 instead. Therefore, we provide the proof for Theorem A.3 in Appendix A.1.
@@
Under Assumptions 3.5 and 3.6, given constant $C$, choose the clipping thresholds $0<C_{1}\leq C_{2}\leq C/B$. Choosing DP noise multiplier $\sigma_{1}$ as
@@
A.1 Proof of Theorem 3.7
Before we provide the convergence proof of Theorem 3.7, let us first define several additional notations which simplify the analysis. We first define a virtual update sequence
$\tilde{\mathbf{x}}^{t}=\mathbf{x}^{t}-\eta\mathbf{e}^{t},$ (8)
From the update rules in Algorithm 2, we can obtain the update rule for the virtual sequence as
$\tilde{\mathbf{x}}^{t+1}$ $=\mathbf{x}^{t+1}-\eta\mathbf{e}^{t+1}$ (9)
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbf{x}^{t}-\eta\mathbf{v}^{t}-\eta\mathbf{w}^{t}-\eta\mathbf{e}^{t+1}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbf{x}^{t}-\eta\mathbf{v}^{t}-\eta\mathbf{w}^{t}-\eta(\mathbf{e}^{t}+\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}\mathbf{g}^{t}_{i}-\mathbf{v}^{t})$
$=\mathbf{x}^{t}-\eta\mathbf{e}^{t}-\frac{\eta}{B}\sum_{i\in\mathcal{B}^{t}}\mathbf{g}^{t}_{i}-\eta\mathbf{w}^{t}$
$=\tilde{\mathbf{x}}^{t}-\frac{\eta}{B}\sum_{i\in\mathcal{B}^{t}}\mathbf{g}^{t}_{i}-\eta\mathbf{w}^{t},$
where $(a)$ applies the update rule of $\mathbf{x}^{t}$ (Algorithm 2 Line 5-6); $(b)$ applies the update rule of $\mathbf{e}^{t}$ (Algorithm 2 Line 7). Notice that the virtual sequence has the same update rule of DPSGD without the clipping operation.
Additionally, let us denote the clipping factors as $\alpha^{t}_{i}:=\min\left\{1,\frac{C_{1}}{\left\lVert\mathbf{g}^{t}_{i}\right\rVert}\right\}$, $\alpha^{t}_{\mathbf{e}}:=\min\left\{1,\frac{C_{2}}{\left\lVert\mathbf{e}^{t}\right\rVert}\right\},$ so that
@@
First, with Assumption 3.2 and let $\operatorname{\mathbb{E}}_{t}$ denote the conditional expectation given history information from iteration 00 up to $t-1$ and condition on the randomness at iteration $t$, then we have:
$\operatorname{\mathbb{E}}_{t}$ $[f(\tilde{\mathbf{x}}^{t+1})-f^{\star}]\leq f(\tilde{\mathbf{x}}^{t})-f^{\star}+\operatorname{\mathbb{E}}_{t}\left\langle\nabla f(\tilde{\mathbf{x}}^{t}),\tilde{\mathbf{x}}^{t+1}-\tilde{\mathbf{x}}^{t}\right\rangle+\operatorname{\mathbb{E}}_{t}\left[\frac{L}{2}\left\lVert\tilde{\mathbf{x}}^{t+1}-\tilde{\mathbf{x}}^{t}\right\rVert^{2}\right]$
=(9)${{=}}f(\tilde{\mathbf{x}}^{t})-f^{\star}+\frac{L\eta^{2}}{2}\operatorname{\mathbb{E}}_{t}\left[\left\lVert\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}\mathbf{g}^{t}_{i}-\mathbf{w}^{t}\right\rVert^{2}\right]-\eta\left\langle\nabla f(\tilde{\mathbf{x}}^{t}),\operatorname{\mathbb{E}}_{t}\left[\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}\mathbf{g}^{t}_{i}-\mathbf{w}^{t}\right]\right\rangle$
$\stackrel{{\scriptstyle(a)}}{{\leq}}f(\tilde{\mathbf{x}}^{t})-f^{\star}-\eta\left\langle\nabla f(\tilde{\mathbf{x}}^{t}),\nabla f(\mathbf{x}^{t})\right\rangle+\frac{L\eta^{2}}{2}\left(\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}+\frac{\sigma^{2}}{B}+d\sigma^{2}_{1}\right)$
=(6)${{=}}f(\tilde{\mathbf{x}}^{t})-f^{\star}-\frac{\eta}{2}\left(\left\lVert\nabla f(\tilde{\mathbf{x}}^{t})\right\rVert^{2}+\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}\right)+\frac{\eta}{2}\left\lVert\nabla f(\tilde{\mathbf{x}}^{t})-\nabla f(\mathbf{x}^{t})\right\rVert^{2}$
$\quad+\frac{L\eta^{2}}{2}\left(\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}+\frac{\sigma^{2}}{B}+d\sigma_{1}^{2}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}f(\tilde{\mathbf{x}}^{t})-f^{\star}-\frac{\eta}{2}\left\lVert\nabla f(\tilde{\mathbf{x}}^{t})\right\rVert^{2}-\frac{\eta(1-L\eta)}{2}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}+\frac{\eta L^{2}}{2}\left\lVert\tilde{\mathbf{x}}^{t}-\mathbf{x}^{t}\right\rVert^{2}$
$\quad+\frac{L\eta^{2}}{2}\left(\frac{\sigma^{2}}{B}+d\sigma_{1}^{2}\right)$
=(9)${{=}}f(\tilde{\mathbf{x}}^{t})-f^{\star}-\frac{\eta}{2}\left\lVert\nabla f(\tilde{\mathbf{x}}^{t})\right\rVert^{2}-\frac{\eta(1-L\eta)}{2}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}+\frac{\eta^{3}L^{2}}{2}\left\lVert\mathbf{e}^{t}\right\rVert^{2}+\frac{L\eta^{2}}{2}\left(\frac{\sigma^{2}}{B}+d\sigma_{1}^{2}\right)$ )
$\stackrel{{\scriptstyle(c)}}{{\leq}}(1-\eta\mu)[f(\tilde{\mathbf{x}}^{t})-f^{\star}]-\frac{\eta(1-L\eta)}{2}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}+\frac{\eta^{3}L^{2}}{2}\left\lVert\mathbf{e}^{t}\right\rVert^{2}+\frac{L\eta^{2}}{2}\left(\frac{\sigma^{2}}{B}+d\sigma_{1}^{2}\right),$ (11)
where $(a)$ applies Assumption 3.5 and use the fact that $\mathbf{w}^{t}$ is an independent Gaussian noise from the stochastic gradient; $(b)$ applies Assumption 3.2 to the third term, and $(c)$ applies Assumption 3.4. Next, we bound $\left\lVert\mathbf{e}^{t}\right\rVert^{2}$ from its update rule:
[…]
(5)${{\leq}}(1+\beta^{t-1})(1-\alpha^{t-1}_{\mathbf{e}})^{2}\left\lVert\mathbf{e}^{t-1}\right\rVert^{2}+(1+\frac{1}{\beta^{t-1}})\operatorname{\mathbb{E}}_{t-1}\left\lVert\frac{1}{B}\sum_{i\in\mathcal{B}^{t-1}}(1-\alpha^{t-1}_{i})\mathbf{g}^{t-1}_{i}\right\rVert^{2}$
=$\stackrel{{\scriptstyle(a)}}{{=}}(1+\beta^{t-1})(1-\alpha^{t-1}_{\mathbf{e}})^{2}\left\lVert\mathbf{e}^{t-1}\right\rVert^{2}+(1+\frac{1}{\beta^{t-1}})\left\lVert\operatorname{\mathbb{E}}_{t-1}[(1-\alpha^{t-1}_{i})\mathbf{g}_{i}^{t-1}]\right\rVert^{2}$
$\quad+(1+\frac{1}{\beta^{t-1}})\operatorname{\mathbb{E}}_{t-1}\left\lVert\frac{1}{B}\sum_{i\in\mathcal{B}^{t-1}}\left((1-\alpha^{t-1}_{i})\mathbf{g}^{t-1}_{i}-\operatorname{\mathbb{E}}_{t-1}[(1-\alpha_{i}^{t-1})\mathbf{g}^{t-1}_{i}]\right)\right\rVert^{2}$
=$\stackrel{{\scriptstyle(b)}}{{=}}(1+\beta^{t-1})(1-\alpha^{t-1}_{\mathbf{e}})^{2}\left\lVert\mathbf{e}^{t-1}\right\rVert^{2}+(1+\frac{1}{\beta^{t-1}})\left\lVert\operatorname{\mathbb{E}}_{t-1}[(1-\alpha^{t-1}_{i})\mathbf{g}_{i}^{t-1}]\right\rVert^{2}$
$\quad+(1+\frac{1}{\beta^{t-1}})\frac{1}{B^{2}}\sum_{i\in\mathcal{B}^{t-1}}\operatorname{\mathbb{E}}_{t-1}\left\lVert(1-\alpha^{t-1}_{i})\mathbf{g}^{t-1}_{i}-\operatorname{\mathbb{E}}_{t-1}[(1-\alpha_{i}^{t-1})\mathbf{g}^{t-1}_{i}]\right\rVert^{2}$
$\quad+(1+\frac{1}{\beta^{t-1}})\frac{1}{B^{2}}\sum_{i\neq j\in\mathcal{B}^{t-1}}\operatorname{\mathbb{E}}_{t-1}\left\langle(1-\alpha^{t-1}_{i})\mathbf{g}^{t-1}_{i}-\operatorname{\mathbb{E}}_{t-1}[(1-\alpha_{i}^{t-1})\mathbf{g}^{t-1}_{i}],(1-\alpha^{t-1}_{j})\mathbf{g}^{t-1}_{j}-\operatorname{\mathbb{E}}_{t-1}[(1-\alpha_{i}^{t-1})\mathbf{g}^{t-1}_{i}]\right\rangle$
$\stackrel{{\scriptstyle(c)}}{{\leq}}(1+\beta^{t-1})(1-\alpha^{t-1}_{\mathbf{e}})^{2}\left\lVert\mathbf{e}^{t-1}\right\rVert^{2}+(1+\frac{1}{\beta^{t-1}})\left\lVert\operatorname{\mathbb{E}}_{t-1}\mathbf{g}_{i}^{t-1}-\operatorname{\mathbb{E}}_{t-1}\alpha^{t-1}_{i}\mathbf{g}_{i}^{t-1}\right\rVert^{2}+(1+\frac{1}{\beta^{t-1}})\frac{\sigma^{2}}{B}$
(5)${{\leq}}(1+\beta^{t-1})(1-\alpha^{t-1}_{\mathbf{e}})^{2}\left\lVert\mathbf{e}^{t}\right\rVert^{2}+2(1+\frac{1}{\beta^{t-1}})\left(\left\lVert\operatorname{\mathbb{E}}_{t-1}\mathbf{g}_{i}^{t-1}\right\rVert^{2}+\left\lVert\operatorname{\mathbb{E}}_{t-1}\alpha^{t-1}_{i}\mathbf{g}_{i}^{t-1}\right\rVert^{2}\right)+(1+\frac{1}{\beta^{t-1}})\frac{\sigma^{2}}{B},$
$\stackrel{{\scriptstyle(d)}}{{\leq}}(1+\beta^{t-1})(1-\alpha^{t-1}_{\mathbf{e}})^{2}\left\lVert\mathbf{e}^{t}\right\rVert^{2}+2(1+\frac{1}{\beta^{t-1}})\left(\left\lVert\nabla f(\mathbf{x}^{t-1})\right\rVert^{2}+C^{2}_{1}\right)+(1+\frac{1}{\beta^{t-1}})\frac{\sigma^{2}}{B},$ (12)
where $(a)$ uses the fact that $\operatorname{\mathbb{E}}[X]^{2}=[\operatorname{\mathbb{E}}X]^{2}+\mathrm{Var}(X)$; $(b)$ expands the last term of (a); $(c)$ applies Lemma A.1 to the second last term and applies the fact that samples in $\mathcal{B}$ are independently sampled from $\mathcal{D}$ so the last term equals to 00; and $(d)$ applies Assumption 3.6 and notice the last term is the clipped gradient so $\left\lVert\operatorname{\mathbb{E}}_{t-1}\alpha^{t-1}_{i}\mathbf{g}_{i}^{t-1}\right\rVert^{2}\leq C_{1}^{2}.$ Adding $\frac{\gamma\eta^{3}L^{2}}{2}\times$(12) to (11), we have:
$\operatorname{\mathbb{E}}_{t}$ $\left[f(\tilde{\mathbf{x}}^{t+1})-f^{\star}+\frac{(1-\gamma)\eta^{3}L^{2}}{2}\left\lVert\mathbf{e}^{t}\right\rVert^{2}+\frac{\eta(1-L\eta)}{2}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}\right]\leq\operatorname{\mathbb{E}}_{t}\left[(1-\eta\mu)f(\tilde{\mathbf{x}}^{t})-f^{\star}\right]$
$\quad+\operatorname{\mathbb{E}}_{t}\left[2\gamma\eta^{3}L^{2}(1+\frac{1}{\beta^{t-1}}))\left\lVert\nabla f(\mathbf{x}^{t-1})\right\rVert^{2}+\frac{\gamma\eta^{3}L^{2}(1+\beta^{t-1})(1-\alpha^{t-1}_{\mathbf{e}})^{2}}{2}\left\lVert\mathbf{e}^{t-1}\right\rVert^{2}\right]$
$\quad+\frac{L\eta^{2}}{2}\left(\frac{(1+\gamma\eta L(1+\frac{1}{\beta^{t-1}}))\sigma^{2}}{B}+d\sigma_{1}^{2}+2\gamma\eta L(1+\frac{1}{\beta^{t-1}})C_{1}^{2}\right),$ (13)
Setting the stepsize $\eta$, and constants $\gamma$, $\beta^{t-1}$ as follows
$0<\eta\leq\min\left\{\frac{1}{\mu},\frac{\alpha^{t-1}_{\mathbf{e}}\left(\sqrt{1+16\gamma\frac{1+\alpha^{t-1}_{\mathbf{e}}}{\alpha^{t-1}_{\mathbf{e}}}}-1\right)}{8\gamma(1+\alpha^{t-1}_{\mathbf{e}})L}\right\},\quad\gamma\leq\frac{1}{2+\alpha^{t-1}_{\mathbf{e}}},\quad\beta^{t-1}=\alpha^{t-1}_{\mathbf{e}},$ (14)
we can guarantee that the following relations hold:
$1-L\eta\geq 4\gamma\eta^{2}L^{2}(1+\frac{1}{\beta^{t-1}}),\;1\geq\eta\mu>0,\;\text{and}\;1-\gamma\geq\gamma(1+\beta^{t-1})(1-\alpha^{t-1}_{\mathbf{e}})^{2}.$
Let us define the following quantities:
$\mathcal{L}^{t}$ $:=f(\tilde{\mathbf{x}}^{t})-f^{\star}+\frac{\gamma\eta^{3}L^{2}}{2}\left\lVert\mathbf{e}^{t}\right\rVert^{2}+\frac{\eta(1-L\eta)}{2}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2},$ (15a)
and $\text{and }\quad\kappa$ $:=\min\left\{\eta\mu,1-\frac{\gamma(1+\beta^{t-1})(1-\alpha^{t}_{\mathbf{e}})^{2}}{1-\gamma},1-\frac{4\gamma\eta^{2}L^{2}(1+\frac{1}{\beta^{t-1}})}{1-L\eta}\right\}$ (15b)
Then we can simplify (13) to
$\operatorname{\mathbb{E}}_{t}[\mathcal{L}^{t+1}]$ $\leq(1-\kappa)\mathcal{L}^{t}+\frac{L\eta^{2}}{2}\left(\frac{(1+\gamma\eta L(1+\frac{1}{\beta^{t-1}}))\sigma^{2}}{B}+d\sigma_{1}^{2}+2\gamma\eta L(1+\frac{1}{\beta^{t-1}})C_{1}^{2}\right)$
$=(1-\kappa)\mathcal{L}^{t}+\frac{L\eta^{2}}{2}\Delta^{t},$ (16)
where we have defined $\Delta^{t}:=\frac{(1+\gamma\eta L(1+\frac{1}{\beta^{t-1}}))\sigma^{2}}{B}+d\sigma_{1}^{2}+2\gamma\eta L(1+\frac{1}{\beta^{t-1}})C_{1}^{2}.$ Therefore, by taking the full expectation, we have:
$\operatorname{\mathbb{E}}[\mathcal{L}^{t+1}]$ $\leq(1-\kappa)\operatorname{\mathbb{E}}[\mathcal{L}^{t}]+\frac{\eta^{2}L}{2}\Delta^{t}$ (17)
$\stackrel{{\scriptstyle(a)}}{{\leq}}(1-\kappa)^{t}\mathcal{L}^{0}+\frac{\eta^{2}L}{2}\sum^{t}_{\tau=0}(1-\kappa)^{t-\tau}\Delta^{\tau}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}(1-\kappa)^{t}\mathcal{L}^{0}+\frac{\eta^{2}L}{2\kappa}\max_{\tau}\{\Delta^{\tau}\},$
where $(a)$ recursively expands $\operatorname{\mathbb{E}}[\mathcal{L}^{t}]$ to $\mathcal{L}^{0}$, and $(b)$ applies $\sum^{t}_{\tau=0}(1-\kappa)^{\tau}=\frac{1-(1-\kappa)^{t+1}}{\kappa}\leq\frac{1}{\kappa}.$
In order to derive the final convergence result, we need to determine $\kappa,\eta,$ and $\Delta^{t}$, which all depend on $\alpha^{t}_{\mathbf{e}}$.
@@
Next, we provide a bound for $\alpha^{t}_{\mathbf{e}}$. Let us consider two cases: 1) $\left\lVert\mathbf{e}^{t}\right\rVert\leq C_{2}$ and $\left\lVert\mathbf{e}^{t}\right\rVert>C_{2}.$ For case 1), it is clear that $\alpha^{t}_{\mathbf{e}}=1.$ For case 2), we have $\alpha^{t}_{\mathbf{e}}=\frac{C_{2}}{\left\lVert\mathbf{e}^{t}\right\rVert}.$ We can further bound $\left\lVert\mathbf{e}^{t}\right\rVert$ recursively as

### v2

@@ 3.2 Theoretical analysis
Assume the problem satisfies Assumptions 3.4, and 3.5, given constant $C$, by fixing the clipping thresholds $0<C_{1}\leq C_{2}\leq C/B$, independent of $G,\sigma$, and assume $\frac{B}{N}\leq\frac{1}{5}$. Choose DP noise standard deviation $\sigma_{1}$ as
@@
Note that although Assumptions 3.4, and 3.5 are used in the proof, the result does not rely on the specific values of the bounds, which can be arbitrarily large. Due to the accumulated influence of the update of $\mathbf{e}^{t}$, the DiceSGD requires larger DP-noise than the DPSGD algorithm (larger by a constant multiplicative factor). The detailed proof is given in Appendix A.2. By optimizing $T$ we have the following utility-privacy trade-off for DiceSGD.
@@
Under the same assumptions of Theorem 3.6, choose the stepsize $\eta=\mathcal{O}(\frac{1}{\sqrt{T}}),$ and clipping thresholds $0<3C_{1}<C_{2}\leq C/B$, and choose noise multiplier $\sigma_{1}^{2}$ as Theorem 3.7. By running DiceSGD for $T=\mathcal{O}\left(\frac{N^{2}\epsilon^{2}}{\tilde{G}\log(1/\delta)}\right)$ iterations, the algorithm guarantees $(\epsilon,\delta)$-DP, while converging to a solution where the loss function satisfies:
$\operatorname{\mathbb{E}}[\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}]=\mathcal{O}\left(\frac{\sqrt{\tilde{G}\log(1/\delta)}}{N\epsilon}\right).$
The corollary indicates that when $N\rightarrow\infty$, the expected loss converges with rate $\mathcal{O}(\frac{\log(N)}{N^{2}})$ with arbitrary clipping thresholds $C_{2}\geq C_{1}>0$ and eliminates the constant clipping bias in DPSGD-GC.
@@ 4 Numerical experiments
In the experiment, we use the similar Adam variant of DPSGD-GC developed following Bu et al. (2021) to implement both DPSGD-GC and DiceSGD (see Appendix C.3 for details). We perform extensive evaluations of DiceSGD on image classification, and natural language processing (NLP) tasks to demonstrate its advantage over DPSGD-GC. The experiments were run on an Intel Xeon W-2102 CPU with an NVIDIA TITAN X GPU for image classification, and on an NVIDIA A100 GPU for NLP tasks. We conduct extra ablation studies on the choice of the clipping threshold $C_{1},C_{2}$ and learning rate $\eta$ on Cifar-10 and Cifar-100 datasets, which show that DiceSGD benefits from using a smaller clipping threshold and choosing $C_{2}=C_{1}$ gives the best result in most cases. More results and discussions are given in Appendix C.1 due to the space limitation.
Image classification. We use both Cifar-10 and Cifar-100 datasets for experiments and use ViT-small (Dosovitskiy et al., 2020) as the training model, which is pre-trained on Imagenet. We fine-tune the model for $3$ epochs with batch size $B=1000$. The stepsize for DPSGD-GC and DiceSGD are selected through grid search from $\eta\in\{10^{-2},10^{-3},10^{-4}\}.$ The experiment results are shown in Table 2.
Table 2: Test accuracy of DPSGD-GC and DiceSGD on Cifar-10 and Cifar-100 datasets with different clipping thresholds and $(2,10^{-5})$-DP.
@@
Natural language processing. To validate the ability of DiceSGD for training larger models on different tasks, we further conduct experiments on the NLP task. Specifically, we fine-tune the GPT-2 model (Radford et al., 2018) on the E2E NLG Challenge for $10$ epochs with batch size $B=1000$, and report the standard metrics such as BLUE, ROUGE-L, etc., used in Hu et al. (2021) for evaluation. The results in Table 3 show that DiceSGD has better performance than DPSGD-GC.
@@ 5 Conclusion
Acknowledgement
M. Hong and X. Zhang are supported by NSF grants CCF 1910385 and EPCN 2311007. X. Zhang is supported by Doctoral Dissertation Fellowship 2023, University of Minnesota.
@@ Appendix A Proof of results in Section 3
In this section, we provide detailed proofs of the theorems in Section 3. We find the following relations useful:
@@
Assume the problem satisfies Assumption 3.1, 3.2, 3.4, and 3.5. Given any constant DP noise multiplier $\sigma_{1}$, by running DiceSGD (Algorithm 2) for $T$ iterations, choosing stepsize $\eta=\sqrt{\frac{2(f(\mathbf{x}^{0})-f^{\star})}{TL(2C_{1}^{2}+3C_{2}^{2}+d\sigma_{1}^{2})}}$, clipping thresholds $C_{2}\geq 3C_{1}+\frac{\sigma}{B}>0$. It satisfies
$\operatorname{\mathbb{E}}_{t}\left[\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}\right]$ $\leq 2\sqrt{\frac{2L(f(\mathbf{x}^{0})-f^{\star})(2C_{1}^{2}+3C_{2}^{2}+d\sigma_{1}^{2})}{T}},$ (7)
where the expectation $\operatorname{\mathbb{E}}_{t}$ is taken over $t\in\{0,\dots,T-1\}$, with probability $A_{t}/\sum^{T-1}_{t=0}A_{t}$, with $\{A_{t}\}\in(0,1]$ being a positive sequence defined in (12).
@@
Under Assumptions 3.4 and 3.5, given constant $C$, choose the clipping thresholds $0<C_{1}\leq C_{2}\leq C/B$. Choosing DP noise multiplier $\sigma_{1}$ as
@@
Let us denote the clipping factors as $\alpha^{t}_{i}:=\min\left\{1,\frac{C_{1}}{\left\lVert\mathbf{g}^{t}_{i}\right\rVert}\right\}$, $\alpha^{t}_{\mathbf{e}}:=\min\left\{1,\frac{C_{2}}{\left\lVert\mathbf{e}^{t}\right\rVert}\right\},$ so that
@@
Note that we have the following bound on $\alpha^{t}_{\mathbf{e}}$: Let us consider two cases: 1) $\left\lVert\mathbf{e}^{t}\right\rVert\leq C_{2}$ and $\left\lVert\mathbf{e}^{t}\right\rVert>C_{2}.$ For case 1), it is clear that $\alpha^{t}_{\mathbf{e}}=1.$ For case 2), we have $\alpha^{t}_{\mathbf{e}}=\frac{C_{2}}{\left\lVert\mathbf{e}^{t}\right\rVert}.$ We can further bound $\left\lVert\mathbf{e}^{t}\right\rVert$ recursively as